189-346/377B: Number Theory Assignment 7

Due: Monday, April 2

1. Section 9.2., Problem 1.

2. Section 9.2., Problem 3.

3. Section 9.3., Problem 1.

4. Section 9.3., Problem 4.

5. 377: Read section 9.1., particularly Theorems 9.1 and 9.2. Show that the p-adic number

p + p^{2} + p^{6} + p^{24} + ⋯ + p^{n!} + ⋯

is transcendental.